OpenStudy (anonymous):

rewrite the quadratic equation -x^2-2x+8 in standard form

OpenStudy (anonymous):

what does standard from means ??!!

OpenStudy (anonymous):

if it's an eqn, I would assume it's equal to zero...so, you can rewrite it as x^2+2x-8=0 complete the square (b/2)^2 = (2/2)^2 = 1 so you want the last term to be 1, so x^2 + 2x + 1 + 7 =0 factor x^2 +2x+1 (leave 7 alone) (x+1)^2 + 7

OpenStudy (anonymous):

BME_21 I tried doing what you did but the (x+1)^2 did not look right.so thank you:)

OpenStudy (anonymous):

You were close...it was (x+1)^2+9 that is what my homework told me